# Federated Continual 3D Segmentation With Single-round Communication

### v1

@@ 1 Introduction
Specifically, we address federated continual learning (FCL) in the context of multi-class 3D segmentation, both under-explored topics compared to fixed FL and to FL-based classification, respectively. Although FL has rapidly evolved, the client-server communication strategy introduced by the classic FedAvg method (McMahan et al., 2017), which performs model aggregation per communication round (MAPCR), still dominates the design of FL methods. However, MAPCR demands frequent client-server communication, strict synchronization, and identical model architectures. In FCL settings, where new clients may join or existing ones may be updated, MAPCR can become impractical. MAPCR requires all clients to have continuous access to all their data, which is often impossible in real-world scenarios due to privacy, licensing, or data storage constraints. Even when data availability is not an issue, MAPCR requires all clients to retrain the updated global model from scratch, significantly increasing both computational and communication overhead.
@@ 2 Related Work
FAug (Jeong et al., 2018) Cls NIID, CE $\bar{z}$ $\bar{z}$ $+$ generator regularizer $+$ generator generator
FedDF (Lin et al., 2020) Cls NIID, MH $\omega$ $\omega$ digestion $+$ $D_{p}$
CFD (Sattler et al., 2021) Cls MH, CE $\tilde{y}_{p}$ $\tilde{y}_{p}$ digestion $+$ $D_{p}$ digestion $+$ $D_{p}$
FedAD (Gong et al., 2021) Cls MH, CE $z_{p},A_{p}$ digestion $+$ $D_{p}$
FedGKT (He et al., 2020) Cls MH, CE $z,H,y$ $z$ regularizer regularizer
Chen et al.  (Chen et al., 2023) Cls P $\omega$ $\omega$ regularizer -
Target (Zhang et al., 2023) Cls CL $\omega$ $\omega+D_{s}$ regularizer generator
Re-Fed (Li et al., 2024) Cls CL $\omega$ $\omega$
CFeD (Ma et al., 2022) Cls NIID, CL $\omega$ $\omega$ digestion $+$ $D_{p}$ digestion + $D_{p}$
Fed-MENU (Xu et al., 2023) Seg NIID $\omega$ $\omega$
Kim et al.  (Kim et al., 2024) Seg NIID $\omega$ $\omega$ regularizer -
Ours Seg NIID, P, CE, MH, CL $\omega$ $\omega$ digestion + $D_{p}$
Table 1: Comparison of related FL methods with ours. FL challenges addressed: data non-iidness (NIID), model heterogeneity (MH), communication efficiency (CE), personalization (P), and continual learning (CL). Symbols: $\omega$ model parameters, $z$ logit vectors (model output before softmax), $\bar{z}$ class-wise average logit vectors, $y$ labels of local data, $\tilde{y}_{p}$ soft targets (model output after softmax) on public data, $H$ intermediate feature maps, $A$ attention maps, $D_{p}$ extra public data, $D_{s}$ synthetic data. Regularizer: uses KD to regularize training. Digestion: uses KD to supervise training.
@@ 2.1 Federated Continual Learning (FCL)
Most current FL frameworks consider a fixed setting with constant training objectives. This traditional FL framework primarily addresses non-IID static data with spatial heterogeneity. Recently, researchers have begun to explore temporal heterogeneity by introducing continual learning within FL (Usmanova et al., 2021; Dong et al., 2022; Ma et al., 2022; Qi et al., 2023; Tran et al., 2024; Babakniya et al., 2024; Zhang et al., 2023; Li et al., 2024). In FCL, the federation updates the global model to adapt to changing client needs. For instance, new clients may join, and existing clients may modify the model to accommodate changes in their objects of interest. Some approaches tackle this challenge using synthetic data replay (Dong et al., 2022; Qi et al., 2023; Tran et al., 2024; Babakniya et al., 2024; Zhang et al., 2023), while others employ memory replay, selectively retaining representative old data locally (Usmanova et al., 2021; Li et al., 2024). In terms of application tasks, most current FCL research has focused on image classification. To the best of our knowledge, our work is the first to investigate FCL in the context of multi-class 3D segmentation.
@@ 2.2 Federated Medical Image Segmentation
Current literature addresses various aspects of federated medical image segmentation; distribution shift (Liu et al., 2021; Xu et al., 2022; Wang et al., 2022, 2023), weak annotations (Zhu et al., 2023; Wicaksana et al., 2022), and noisy labels (Wu et al., 2024). Most studies focus on binary segmentation, while multi-class segmentation remains relatively underexplored, despite its practical importance. This paper considers federated multi-class segmentation for 3D abdominal CT scans. For this task, Xu et al. proposed class-specific encoders to improve feature extraction (Xu et al., 2023). Kim et al. used knowledge distillation for regularization (Kim et al., 2024). As summarized in Table 1, both methods rely on frequent client-server communication and are limited to fixed FL scenarios. In contrast, our approach is designed for FCL scenarios and requires only one-time model aggregation to generate the global model.
@@ 2.3 Knowledge Distillation (KD) in FL
KD (Hinton, 2015) has been widely adopted in FL due to its ability to transfer knowledge without requiring access to the original training data. For example, KD can be used to improve communication efficiency (CE) (Jeong et al., 2018), to alleviate model heterogeneity (MH) (Lin et al., 2020; Afonin and Karimireddy, 2021) and to support personalization (P) (Chen et al., 2023). Additionally, KD has been applied to mitigate forgetting in FCL (Ma et al., 2022; Zhang et al., 2023; Li et al., 2024). Table 1 compares our approach with representative, and closely related KD-based FL methods  [^1] For a survey covering a broader scope, refer to Mora et al. (2022).. Existing methods often address only one or two specific FL challenges, such as CE, MH, P, or FCL. However, in real-world FL scenarios, these challenges are often interconnected, which requires a unified solution to address them simultaneously. Furthermore, most methods rely on the MAPCR strategy, which requires frequent client-server communication. We argue that the frequent communication of MAPCR imposes limitations in tackling challenges such as communication synchronization and FCL. To address these interconnected challenges, we propose a novel framework that does not use MAPCR while offering a unified solution to challenges such as CE, MH, P, and FCL.
@@ 4 Methodology
12:    $\tilde{y}\leftarrow\text{PredictionAggregation}(\hat{y}^{1},\hat{y}^{2},...,\hat{y}^{|S_{t}|})$triangleright Equation 1
13:    $w^{G}_{t}\leftarrow\text{MultiModelDistillation}(\mathcal{D}_{dist},\tilde{y},w_{0}^{G})$triangleright Equations 2, 3, 4
@@ 4.1 Local Model Training
In our framework, each client first trains its model locally using annotated data through a standard supervised learning pipeline. Since clients may annotate images for different purposes, they often focus on different segmentation targets, leading to variations in the number of prediction classes and class overlaps across clients. To accommodate mismatched label space across clients, we employ CLIP-DrivenSeg (Liu et al., 2023) as our backbone. This network leverages CLIP embeddings (Radford et al., 2021) to represent each class, allowing the network to dynamically extend its prediction classes without structural modifications. During local supervised training, as annotations are limited to a subset of target classes relevant to each client, binary cross-entropy loss and binary dice loss are used. These losses are computed and back-propagated exclusively for the classes with corresponding annotations, ensuring that precise supervision is applied only to the labeled classes.
@@ 4.2 Server Distillation
After local training, client models are sent to the server for aggregation. On the server side, knowledge distillation (KD) (Hinton, 2015) is utilized for model aggregation, using a publicly available unlabeled dataset as the distillation dataset. Pseudo-labels for the distillation dataset are generated based on predictions from the client models. Each client runs a one-time inference on the distillation dataset to produce its predictions, which are then aggregated to create the pseudo-labels. Given that clients may have overlapping segmentation targets, an effective pseudo-label aggregation strategy is required. For classes with predictions from multiple models, entropy impurity (Shen et al., 2019) is used to quantify prediction ambiguity. For each distillation sample, the prediction with the lowest entropy impurity, indicating the highest confidence, is selected as the pseudo-label for that class. The entropy impurity for class $c$ for a given input image is calculated as follows:
@@
Figure 3: Comparison of communication and computation loads between the classic MAPCR-type method FedAvg and the proposed method during 5-stage FCL segmentation across five clients. Here 1×1$\times$1 × represents the computation required to train a single model to convergence (assumed to be similar across client models).
@@ 4.3 Communication and Computation Complexity Analysis
Figure 7: Out-of-distribution DICE score and ASSD performance of the global model after 5-stage FCL, evaluated on the BTCV dataset (Landman et al., 2015).
@@ 5.1 Dataset
We performed experiments with six publicly available 3D abdominal CT segmentation datasets, encompassing annotations for 19 abdominal organs and associated pathologies. These datasets are: 1) LiTS dataset (Bilic et al., 2023); 2) KiTS dataset (Heller et al., 2019); 3) MSD pancreas dataset (Antonelli et al., 2022); 4) MSD spleen dataset (Antonelli et al., 2022); 5) AMOS dataset (Ji et al., 2022); 6) BTCV dataset (Landman et al., 2015). Additionally, the WORD dataset (Luo et al., 2021) is utilized as the distillation dataset. For the detailed information of all these datasets, please refer to supplementary material A.
@@ 5.3 Task Setup
To replicate real-world multi-site scenarios, we treat each dataset as a distinct client and utilize all available annotations from it. This allows for overlap in objects of interest across clients, while also introducing unique, client-specific objects. As we focus on FCL in this work, clients are incrementally introduced to the federation. The process begins with two clients (stage 1): MSD-pancreas (Antonelli et al., 2022) and MSD-spleen (Antonelli et al., 2022). Subsequently, KiTS (Heller et al., 2019), LiTS (Bilic et al., 2023), and AMOS (Ji et al., 2022) are added. To simulate evolving clients, the AMOS dataset, which includes annotations for multiple organs, is divided into two sets with distinct classes. This setup establishes a five-stage FCL scenario, where each stage introduces a new client (stage 2, 3, 4) or updates an existing one (stage 5).
@@ 5.4 Implementation Details
The proposed method is implemented in PyTorch and CLIP-DrivenSeg (Liu et al., 2023) is used as the segmentation backbone. We follow its original training strategy utilizing the AdamW optimizer with a warm-up scheduler over 50 epochs. The initial learning rate is set to $4e^{-4}$, with a momentum of 0.9 and a weight decay of $1e^{-5}$. For methods involving MAPCR, the model is trained for 1000 communication rounds, with 1 epoch per round. For methods without MAPCR, the model undergoes training for 1000 epochs.
@@ 5.5 Comparison with Benchmarks
We also compare the segmentation performance of our proposed method with centralized training. Centralized training, which aggregates data from all clients for model development, serves as the performance upper bound. Figure 4 presents the results. After five stages of FCL, the proposed MMDS method achieves an average DICE score of 78.65% across all organ classes, compared to 85.09% for centralized training. In 12 out of 15 organ classes, MMDS achieves a DICE score greater than 70%. However, its accuracy decreases for challenging small or elongated structures like the pancreas and the duodenum. On the other hand, the performance of pathology classes, such as tumors, is notably lower. We conjecture this is because the distillation dataset, WORD (Luo et al., 2021), does not include pathology examples, limiting the activation of pathological features during distillation. Consequently, the global model lacks sufficient information to effectively learn these classes. Addressing the challenge of distilling knowledge from the teacher model without heavy reliance on distillation data remains an open question in knowledge distillation. We plan to explore this issue in future work.
@@ 5.7 Heterogeneous Model Architecture
Our approach supports model agnosticism in FL by allowing each client model to be independently and fully trained. The global model is then trained by aligning the input-output functions of the client models eliminating the need for homogeneous model architectures. To demonstrate the flexibility of the proposed method in handling heterogeneous model structures, we employ two types of backbones: convolutional-based and transformer-based 3D U-Nets. Specifically, Swin UNETR (Hatamizadeh et al., 2021) is used as the baseline feature extractor for the second and fourth clients, while 3D U-Net (Çiçek et al., 2016) is employed for the remaining clients. The five-step FCL process is repeated. Figure 6 presents the performance of the final global model. The results demonstrate that the proposed method effectively aggregates client models with diverse architectures, achieving performance comparable to setups where clients use homogeneous model structures.
@@ 5.8 Out-of-Distribution Performance
To assess the robustness of the proposed method, we evaluate its out-of-distribution performance using the BTCV dataset (Landman et al., 2015), which represents an external client not involved in the federation. Figure 7 presents the results. Since the ”Portal Vein and Splenic Vein” class is not included in the federation (as none of the participating clients contained this class), it is excluded from the evaluation. For comparison, the performance of the centrally trained model is provided as the upper bound. As shown in the figure, the proposed method achieves comparable performance to the upper bound on most classes, demonstrating its robustness.
@@ Appendix A Dataset
In this section, we introduce the 3D abdominal CT segmentation datasets used in our experiments. Since different datasets provide annotations for varying classes, with some overlap between them, we followed the approach outlined in Liu et al. (2023) to unify the label indices across all datasets. Overall, we focus on segmenting 19 abdominal organs and pathological structures. As many datasets do not provide ground truth annotations for their test data, following Liu et al. (2023), we reorganize the available training sets, splitting them into training, validation, and testing subsets. The detailed information for each dataset is provided below, and a summary is presented in Table 3, including the number of training samples, the number of class labels, and the label lists.
@@
BTCV(Landman et al., 2015) comprises 30 scans with volume size range from $[85\sim 198]\times[512]\times[512]$ pixels in dimensions $D\times H\times W$, , where $D$ represents the depth, $H$ is the height, and $W$ is the width. As the BTCV dataset is used for out-of-distribution performance evaluation, all its data are used for testing. The in-plane pixel spacing in this dataset varies between 0.59mm to 0.98mm, and the slice thickness ranges from 2.50mm to 5.00mm. The dataset provides pixel-wise annotations for 14 classes: spleen, right Kidney, left kidney, gall bladder, esophagus, liver, stomach, aorta, postcava, portal vein and splenic vein, pancreas, right adrenal gland, left adrenal gland, and duodenum.
@@
LiTS (2017) (Bilic et al., 2023) comprises 131 scans with volume size range from $[74\sim 987]\times[512]\times[512]$ pixels in dimensions $D\times H\times W$. Out of the 131 scans, 95 are used for training, 16 for validation, and 20 for testing. The in-plane pixel spacing in this dataset varies between 0.56mm and 1.00mm, and the slice thickness ranges from 0.70mm to 5.00mm. The dataset provides annotations for 2 classes: liver and liver tumor.
@@
KiTS (Heller et al., 2019) comprises 210 scans with volume sizes ranging from $[29\sim 1059]\times[512]\times[512,796]$ pixels in dimensions $D\times H\times W$. Out of the 210 scans, 142 are used for training, 22 for validation, and 46 for testing. The in-plane pixel spacing in this dataset varies between 0.44mm and 1.04mm, and the slice thickness ranges from 0.50mm to 5.00mm. The dataset provides annotations for 4 classes: left kidney, right kidney, kidney tumor, and kidney crystal.
@@
MSD - Spleen (Antonelli et al., 2022) comprises 41 scans with volume size range from $[31\sim 168]\times[512]\times[512]$ pixels in dimensions $D\times H\times W$. Out of the 41 scans, 27 are used for training, 4 for validation, and 10 for testing. The in-plane pixel spacing in this dataset varies between 0.61mm and 0.98mm, and the slice thickness ranges from 1.50mm to 8.00mm. The dataset provides annotations for 1 class: spleen.
@@
MSD - Pancreas (Antonelli et al., 2022) comprises 281 scans with volume size range from $[37\sim 751]\times[512]\times[512]$ pixels in dimensions $D\times H\times W$. Out of the 281 scans, 198 are used for training, 29 for validation, and 54 for testing. The in-plane pixel spacing in this dataset varies between 0.54mm and 0.98mm, and the slice thickness ranges from 0.63mm to 7.50mm. The dataset provides annotations for 2 classes: pancreas and pancreas tumor.
@@
AMOS (Ji et al., 2022) comprises 200 scans with volume size range from $[64\sim 512]\times[60~{}768]\times[192\sim 768]$ pixels in dimensions $D\times H\times W$. Out of the 200 scans, 145 are used for training, 23 for validation, and 32 for testing. The in-plane pixel spacing in this dataset varies between 0.45mm and 3.00mm, and the slice thickness ranges from 0.82mm to 5.00mm. The dataset provides annotations for 15 classes: spleen, right kidney, left kidney, gall bladder, esophagus, liver, stomach, aorta, postcava, pancreas, right adrenal gland, left adrenal gland, duodenum, bladder, and prostate.

### v2

@@ 1 Introduction
Specifically, we address federated continual learning (FCL) in the context of multi-class 3D segmentation, both under-explored topics compared to fixed FL and to FL-based classification, respectively. Although FL has rapidly evolved, the client-server communication strategy introduced by the classic FedAvg method (mcmahan2017communication), which performs model aggregation per communication round (MAPCR), still dominates the design of FL methods. However, MAPCR demands frequent client-server communication, strict synchronization, and identical model architectures. In FCL settings, where new clients may join or existing ones may be updated, MAPCR can become impractical. MAPCR requires all clients to have continuous access to all their data, which is often impossible in real-world scenarios due to privacy, licensing, or data storage constraints. Even when data availability is not an issue, MAPCR requires all clients to retrain the updated global model from scratch, significantly increasing both computational and communication overhead.
@@ 2 Related Work
FAug (jeong2018communication) Cls NIID, CE $\bar{z}$ $\bar{z}$ ++ generator regularizer ++ generator generator
FedDF (lin2020ensemble) Cls NIID, MH $\omega$ $\omega$ - digestion ++ $D_{p}$
CFD (sattler2021cfd) Cls MH, CE $\tilde{y}_{p}$ $\tilde{y}_{p}$ digestion ++ $D_{p}$ digestion ++ $D_{p}$
FedAD (gong2021ensemble) Cls MH, CE $z_{p},A_{p}$ - - digestion ++ $D_{p}$
FedGKT (he2020group) Cls MH, CE $z,H,y$ $z$ regularizer regularizer
Chen et al.  (chen2023spectral) Cls P $\omega$ $\omega$ regularizer -
Target (zhang2023target) Cls CL $\omega$ $\omega+D_{s}$ regularizer generator
Re-Fed (li2024towards) Cls CL $\omega$ $\omega$ - -
CFeD (ma2022continual) Cls NIID, CL $\omega$ $\omega$ digestion ++ $D_{p}$ digestion + $D_{p}$
Fed-MENU (xu2023federated) Seg NIID $\omega$ $\omega$ - -
Kim et al.  (kim2024federated) Seg NIID $\omega$ $\omega$ regularizer -
Ours Seg NIID, P, CE, MH, CL $\omega$ $\omega$ - digestion + $D_{p}$
Table 1: Comparison of related FL methods with ours. FL challenges addressed: data non-iidness (NIID), model heterogeneity (MH), communication efficiency (CE), personalization (P), and continual learning (CL). Symbols: $\omega$ model parameters, $z$ logit vectors (model output before softmax), $\bar{z}$ class-wise average logit vectors, $y$ labels of local data, $\tilde{y}_{p}$ soft targets (model output after softmax) on public data, $H$ intermediate feature maps, AA attention maps, $D_{p}$ extra public data, $D_{s}$ synthetic data. Regularizer: uses KD to regularize training. Digestion: uses KD to supervise training.
@@ 2.1 Federated Continual Learning (FCL)
Most current FL frameworks consider a fixed setting with constant training objectives. This traditional FL framework primarily addresses non-IID static data with spatial heterogeneity. Recently, researchers have begun to explore temporal heterogeneity by introducing continual learning within FL (usmanova2021distillation; dong2022federated; ma2022continual; qi2023better; tran2024text; babakniya2024data; zhang2023target; li2024towards). In FCL, the federation updates the global model to adapt to changing client needs. For instance, new clients may join, and existing clients may modify the model to accommodate changes in their objects of interest. Some approaches tackle this challenge using synthetic data replay (dong2022federated; qi2023better; tran2024text; babakniya2024data; zhang2023target), while others employ memory replay, selectively retaining representative old data locally (usmanova2021distillation; li2024towards). In terms of application tasks, most current FCL research has focused on image classification. To the best of our knowledge, our work is the first to investigate FCL in the context of multi-class 3D segmentation.
@@ 2.2 Federated Medical Image Segmentation
Current literature addresses various aspects of federated medical image segmentation; distribution shift (liu2021feddg; xu2022closing; wang2022personalizing; wang2023feddp), weak annotations (zhu2023feddm; wicaksana2022fedmix), and noisy labels (wu2024feda3i). Most studies focus on binary segmentation, while multi-class segmentation remains relatively underexplored, despite its practical importance. This paper considers federated multi-class segmentation for 3D abdominal CT scans. For this task, Xu et al. proposed class-specific encoders to improve feature extraction (xu2023federated). Kim et al. used knowledge distillation for regularization (kim2024federated). As summarized in Table 1, both methods rely on frequent client-server communication and are limited to fixed FL scenarios. In contrast, our approach is designed for FCL scenarios and requires only one-time model aggregation to generate the global model.
@@ 2.3 Knowledge Distillation (KD) in FL
KD (hinton2015distilling) has been widely adopted in FL due to its ability to transfer knowledge without requiring access to the original training data. For example, KD can be used to improve communication efficiency (CE) (jeong2018communication), to alleviate model heterogeneity (MH) (lin2020ensemble; afonin2021towards) and to support personalization (P) (chen2023spectral). Additionally, KD has been applied to mitigate forgetting in FCL (ma2022continual; zhang2023target; li2024towards). Table 1 compares our approach with representative, and closely related KD-based FL methods  [^1] For a survey covering a broader scope, refer to mora2024knowledge.. Existing methods often address only one or two specific FL challenges, such as CE, MH, P, or FCL. However, in real-world FL scenarios, these challenges are often interconnected, which requires a unified solution to address them simultaneously. Furthermore, most methods rely on the MAPCR strategy, which requires frequent client-server communication. We argue that the frequent communication of MAPCR imposes limitations in tackling challenges such as communication synchronization and FCL. To address these interconnected challenges, we propose a novel framework that does not use MAPCR while offering a unified solution to challenges such as CE, MH, P, and FCL.
@@ 4 Methodology
12:  $\tilde{y}\leftarrow\text{PredictionAggregation}(\hat{y}^{1},\hat{y}^{2},...,\hat{y}^{|S_{t}|})$ $\triangleright$ Equation 1
13:  $w^{G}_{t}\leftarrow\text{MultiModelDistillation}(\mathcal{D}_{dist},\tilde{y},w_{0}^{G})$ $\triangleright$ Equations 2, 3, 4
@@ 4.1 Local Model Training
In our framework, each client first trains its model locally using annotated data through a standard supervised learning pipeline. Since clients may annotate images for different purposes, they often focus on different segmentation targets, leading to variations in the number of prediction classes and class overlaps across clients. To accommodate mismatched label space across clients, we employ CLIP-DrivenSeg (liu2023clip) as our backbone. This network leverages CLIP embeddings (radford2021learning) to represent each class, allowing the network to dynamically extend its prediction classes without structural modifications. During local supervised training, as annotations are limited to a subset of target classes relevant to each client, binary cross-entropy loss and binary dice loss are used. These losses are computed and back-propagated exclusively for the classes with corresponding annotations, ensuring that precise supervision is applied only to the labeled classes.
@@ 4.2 Server Distillation
After local training, client models are sent to the server for aggregation. On the server side, knowledge distillation (KD) (hinton2015distilling) is utilized for model aggregation, using a publicly available unlabeled dataset as the distillation dataset. Pseudo-labels for the distillation dataset are generated based on predictions from the client models. Each client runs a one-time inference on the distillation dataset to produce its predictions, which are then aggregated to create the pseudo-labels. Given that clients may have overlapping segmentation targets, an effective pseudo-label aggregation strategy is required. For classes with predictions from multiple models, entropy impurity (shen2019customizing) is used to quantify prediction ambiguity. For each distillation sample, the prediction with the lowest entropy impurity, indicating the highest confidence, is selected as the pseudo-label for that class. The entropy impurity for class $c$ for a given input image is calculated as follows:
@@
Figure 3: Comparison of communication and computation loads between the classic MAPCR-type method FedAvg and the proposed method during 5-stage FCL segmentation across five clients. Here $1\times$ represents the computation required to train a single model to convergence (assumed to be similar across client models).
@@ 4.3 Communication and Computation Complexity Analysis
Figure 7: Out-of-distribution DICE score and ASSD performance of the global model after 5-stage FCL, evaluated on the BTCV dataset (landman2015miccai).
@@ 5.1 Dataset
We performed experiments with six publicly available 3D abdominal CT segmentation datasets, encompassing annotations for 19 abdominal organs and associated pathologies. These datasets are: 1) LiTS dataset (bilic2023liver); 2) KiTS dataset (heller2019kits19); 3) MSD pancreas dataset (antonelli2022medical); 4) MSD spleen dataset (antonelli2022medical); 5) AMOS dataset (ji2022amos); 6) BTCV dataset (landman2015miccai). Additionally, the WORD dataset (luo2021word) is utilized as the distillation dataset. For the detailed information of all these datasets, please refer to supplementary material A.
@@ 5.3 Task Setup
To replicate real-world multi-site scenarios, we treat each dataset as a distinct client and utilize all available annotations from it. This allows for overlap in objects of interest across clients, while also introducing unique, client-specific objects. As we focus on FCL in this work, clients are incrementally introduced to the federation. The process begins with two clients (stage 1): MSD-pancreas (antonelli2022medical) and MSD-spleen (antonelli2022medical). Subsequently, KiTS (heller2019kits19), LiTS (bilic2023liver), and AMOS (ji2022amos) are added. To simulate evolving clients, the AMOS dataset, which includes annotations for multiple organs, is divided into two sets with distinct classes. This setup establishes a five-stage FCL scenario, where each stage introduces a new client (stage 2, 3, 4) or updates an existing one (stage 5).
@@ 5.4 Implementation Details
The proposed method is implemented in PyTorch and CLIP-DrivenSeg (liu2023clip) is used as the segmentation backbone. We follow its original training strategy utilizing the AdamW optimizer with a warm-up scheduler over 50 epochs. The initial learning rate is set to $4e^{-4}$, with a momentum of 0.9 and a weight decay of $1e^{-5}$. For methods involving MAPCR, the model is trained for 1000 communication rounds, with 1 epoch per round. For methods without MAPCR, the model undergoes training for 1000 epochs.
@@ 5.5 Comparison with Benchmarks
We also compare the segmentation performance of our proposed method with centralized training. Centralized training, which aggregates data from all clients for model development, serves as the performance upper bound. Figure 4 presents the results. After five stages of FCL, the proposed MMDS method achieves an average DICE score of 78.65% across all organ classes, compared to 85.09% for centralized training. In 12 out of 15 organ classes, MMDS achieves a DICE score greater than 70%. However, its accuracy decreases for challenging small or elongated structures like the pancreas and the duodenum. On the other hand, the performance of pathology classes, such as tumors, is notably lower. We conjecture this is because the distillation dataset, WORD (luo2021word), does not include pathology examples, limiting the activation of pathological features during distillation. Consequently, the global model lacks sufficient information to effectively learn these classes. Addressing the challenge of distilling knowledge from the teacher model without heavy reliance on distillation data remains an open question in knowledge distillation. We plan to explore this issue in future work.
@@ 5.7 Heterogeneous Model Architecture
Our approach supports model agnosticism in FL by allowing each client model to be independently and fully trained. The global model is then trained by aligning the input-output functions of the client models eliminating the need for homogeneous model architectures. To demonstrate the flexibility of the proposed method in handling heterogeneous model structures, we employ two types of backbones: convolutional-based and transformer-based 3D U-Nets. Specifically, Swin UNETR (hatamizadeh2021swin) is used as the baseline feature extractor for the second and fourth clients, while 3D U-Net (cciccek20163d) is employed for the remaining clients. The five-step FCL process is repeated. Figure 6 presents the performance of the final global model. The results demonstrate that the proposed method effectively aggregates client models with diverse architectures, achieving performance comparable to setups where clients use homogeneous model structures.
@@ 5.8 Out-of-Distribution Performance
To assess the robustness of the proposed method, we evaluate its out-of-distribution performance using the BTCV dataset (landman2015miccai), which represents an external client not involved in the federation. Figure 7 presents the results. Since the ”Portal Vein and Splenic Vein” class is not included in the federation (as none of the participating clients contained this class), it is excluded from the evaluation. For comparison, the performance of the centrally trained model is provided as the upper bound. As shown in the figure, the proposed method achieves comparable performance to the upper bound on most classes, demonstrating its robustness.
@@ Appendix A Dataset
In this section, we introduce the 3D abdominal CT segmentation datasets used in our experiments. Since different datasets provide annotations for varying classes, with some overlap between them, we followed the approach outlined in liu2023clip to unify the label indices across all datasets. Overall, we focus on segmenting 19 abdominal organs and pathological structures. As many datasets do not provide ground truth annotations for their test data, following liu2023clip, we reorganize the available training sets, splitting them into training, validation, and testing subsets. The detailed information for each dataset is provided below, and a summary is presented in Table 3, including the number of training samples, the number of class labels, and the label lists.
@@
BTCV(landman2015miccai) comprises 30 scans with volume size range from $85\sim 198]\times[512]\times[512]$ pixels in dimensions $D\times H\times W$, , where $D$ represents the depth, $H$ is the height, and $W$ is the width. As the BTCV dataset is used for out-of-distribution performance evaluation, all its data are used for testing. The in-plane pixel spacing in this dataset varies between 0.59mm to 0.98mm, and the slice thickness ranges from 2.50mm to 5.00mm. The dataset provides pixel-wise annotations for 14 classes: spleen, right Kidney, left kidney, gall bladder, esophagus, liver, stomach, aorta, postcava, portal vein and splenic vein, pancreas, right adrenal gland, left adrenal gland, and duodenum.
@@
LiTS (2017) (bilic2023liver) comprises 131 scans with volume size range from $74\sim 987]\times[512]\times[512]$ pixels in dimensions $D\times H\times W$. Out of the 131 scans, 95 are used for training, 16 for validation, and 20 for testing. The in-plane pixel spacing in this dataset varies between 0.56mm and 1.00mm, and the slice thickness ranges from 0.70mm to 5.00mm. The dataset provides annotations for 2 classes: liver and liver tumor.
@@
KiTS (heller2019kits19) comprises 210 scans with volume sizes ranging from $29\sim 1059]\times[512]\times[512,796]$ pixels in dimensions $D\times H\times W$. Out of the 210 scans, 142 are used for training, 22 for validation, and 46 for testing. The in-plane pixel spacing in this dataset varies between 0.44mm and 1.04mm, and the slice thickness ranges from 0.50mm to 5.00mm. The dataset provides annotations for 4 classes: left kidney, right kidney, kidney tumor, and kidney crystal.
@@
MSD - Spleen (antonelli2022medical) comprises 41 scans with volume size range from $31\sim 168]\times[512]\times[512]$ pixels in dimensions $D\times H\times W$. Out of the 41 scans, 27 are used for training, 4 for validation, and 10 for testing. The in-plane pixel spacing in this dataset varies between 0.61mm and 0.98mm, and the slice thickness ranges from 1.50mm to 8.00mm. The dataset provides annotations for 1 class: spleen.
@@
MSD - Pancreas (antonelli2022medical) comprises 281 scans with volume size range from $37\sim 751]\times[512]\times[512]$ pixels in dimensions $D\times H\times W$. Out of the 281 scans, 198 are used for training, 29 for validation, and 54 for testing. The in-plane pixel spacing in this dataset varies between 0.54mm and 0.98mm, and the slice thickness ranges from 0.63mm to 7.50mm. The dataset provides annotations for 2 classes: pancreas and pancreas tumor.
@@
AMOS (ji2022amos) comprises 200 scans with volume size range from $64\sim 512]\times[60~768]\times[192\sim 768]$ pixels in dimensions $D\times H\times W$. Out of the 200 scans, 145 are used for training, 23 for validation, and 32 for testing. The in-plane pixel spacing in this dataset varies between 0.45mm and 3.00mm, and the slice thickness ranges from 0.82mm to 5.00mm. The dataset provides annotations for 15 classes: spleen, right kidney, left kidney, gall bladder, esophagus, liver, stomach, aorta, postcava, pancreas, right adrenal gland, left adrenal gland, duodenum, bladder, and prostate.